You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100410 (14 Apr 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 April09-10

Observer: A. Fontana, K.Boutsia, R. Speziali
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LUCIFER1

Summary

Roberto S. fixed the shutter of the blue camera. Camera has warmed up during the maintenance in the lab. It has been pumped and cooled.
Good seeing throughout the night (0.7-1.0).
7min lost due to switching on LBC red.

Data Summary

Project Objects Filter ExpTime Notes
Grazian GOODS-N 200H+K OrderSep 3hrs arc -flat - tell
Mannucci Q1422 210zJHK K 1hr arc-flat - NO tell
Mannucci Q1422 210zJHK H 2hr arc- flat- tell

Details

01:00 - Opening the dome - starting interface

01:10 - Narrow band Sky Flats 0001-0085

01:57 - Pointing Star High Elevation (WT10_303) 0086

02:01 - Pointing Star Low Elevation (WT10_097) 0087

02:04 - Pointing Star High Elevation (WT10_303) 0088

02:08 - Pointing Star near the first observing field of the night (GOODS-WT10_385) 0089

02:13 - Persson standard BS9145 for collimation (POSITION,ACTIVE)

02:21 - Collimation - preset to Grazian Point1 acquisition script 0090-0091
  • through slit image 0092-0093

02:47 - Grazian (GOODS) Point1 sci script 0094

02:55 - Abort script - preset was lost due to switching on LBC red (7min lost)
  • aquisition of Grazian (GOODS) field again 0095-0097

03:07 - Repeat Grazian (GOODS) Point1 sci script (3rd hour) 0098-0109
  • N1.8 - 200H+K - OrderSep / 300secx1x12offset (seeing 0.8)

04:15 - Mirror acquisition to see if sources are in slit
  • sources are offset - we perform MOS acquisition again 0110-0115

04:42 - Grazian Point1 sci script (4th hour) 0116-0127

05:50 - Through slit image to see if sources are centered 0128

05:56 - Grazian Point1 sci script (5th hour) 0129-0140

07:02 - Calibration unit ON - Flat script for Grazian
  • Halo2 lamp off 0141-0143
  • Halo2 lamp on 0144-0146

07:06 - Arc acquisition script
  • Dark Ar 0147
  • ARC Ar 0148
  • Dark Ne 0149
  • ARC Ne 0150

07:15 - Telluric star acquisition script
  • Abort - no guide star is found

07:20 - Pointing correction (WT10-385) 0151

07:24 - Telluric star preset (Hipp62402) 0152
  • We calculate offsets for centering the star in the slits

07:33 - Telluric SCI script with the Grazian Mask (913423)
  • the offset turned out to be wrong 0153-0157

07:40 - Telluric ACQ script - repeat 0158

07:43 - Telluric SCI with Grazian Mask (913423) 0159-162
  • We need to perform ARC Mannucci with Grazian Mask for telluric wavelength calibration
  • More images with the mask in turnout 0163-0165 and in FPU 0166

08:04 - Slew to Mannucci Q1422 - MOS acquisition 0167-0170

08:27 - Mannucci Q1422 SCI script K_4 (4th hr) 0171-0182
  • N1.8 - 210_zJHK- K / 300secx1x12offset (seeing 0.8)

09:43 - Mannucci Q1422 Sci script for H band (1st hr-H_1) 0184-0195
  • Through slit image to see if sources are centered 0183

10:55 - Mannucci Q1422 sci script for H_2 0196-0208
  • first images are offset by 1.5 arcsec from aquisition positiion 0196-200
  • correct offset applicated - mask restarted and aborted after 8th exposure 201-208
  • OB considered executed

12:00 - Pointing correction 0209-0210

12:05 - Telluric aquisition in the H band after correcting pointing 0211-0212

12:16 - Telluric SCI script (mannucci_tell_int_Q1422_H.txt) 0213-0220

12:23 - Telluric acquisition in the K band failed- too bright to guide

12:25 - Calibration unit on- Arc acquisition with the Mannucci Mask (933911)
  • ARC K filter 0221-0224
  • ARC H filter 0225-0228

12:40 - Flat script with the Mannucci Mask (933911)
  • Flat H filter - Halo1 OFF 0229 and 0231-0234 and 0237-0242 and 0246
  • Flat H filter - Halo1 ON 0230 and 235-236 and 0243-0245
  • Flat K filter - Halo2 OFF 0247-0251 and 0256
  • Flat K filter - Halo2 ON 0252-0255
  • There is a mistake calculating the lamps time because they turn off before the last 'ON' acquisition

12:55 - ARC and FLAT scripts with Mannucci filter set (210_zJHK/K) and Grazian Mask (913423) for the calibration of yesterday's telluric
  • ARCS Ne OFF 0257
  • ARCS Ne ON 0258
  • ARCS Xe OFF 0259
  • ARCS Xe ON 0258
  • FLAT Halo2 OFF 5sec 0261-0265
  • FLAT Halo2 ON 5sec 0266-0270
  • FLAT Halo2 OFF 25sec 0271-0273-275
  • FLAT Halo2 ON 25sec 0272-0274-0276
  • last ON acquisition is done with the Halo2 lamp off - probably the same error in calculating lamp duration

13:37 - DoDarks.txt script - End of night
  • 30secx2ndit 0277-0282
  • 300secx1ndit 0283-0292

-- VincenzoTest - 10 Apr 2010
Topic revision: r12 - 14 Apr 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback